### v2

@@ 2 Model
Let $i\in\left\{1,...,n\right\}$ index $n$ agents and $t\in\left\{1,...,T\right\}$ index finite $T\geq 2$ periods. Let $d_{X}$ be the dimension of $X_{it}$ and similarly for $Z_{it},$ $A_{i}$, $\varepsilon_{it}$, and $\eta_{it}$. Consider the following baseline TERC model
$Y_{it}$ $=X_{it}^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right),$ (2.1)
@@
where:
  • $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)\in\mathbb{R}^{d_{X}}$, the central object of interest, is a vector of random coefficients modeled as a vector-valued, unknown, and potentially nonlinear function $\beta\left(\cdot\right)$ of $A_{i}\text{ and }\varepsilon_{it}$. Here, $A_{i}$ represents a fixed effect and $\varepsilon_{it}$ drives the time-varying behavior of $\beta_{it}$, both of which can be either scalar- or vector-valued. Furthermore, $A_{i}$ and $\varepsilon_{it}$ can be arbitrarily correlated.
  • $\eta_{it}\in\mathbb{R}^{d_{\eta}}$ is a continuously distributed and time-varying random vector that influences the value of $X_{it}$. $\eta_{it}$ is allowed to be correlated with both $A_{i}$ and $\varepsilon_{it}$. Additionally, different coordinates of $\eta_{it}$ can also be correlated.
  • $X_{it}\in\mathbb{R}^{d_{X}}$ is a vector of choice variables for agent $i$ in period $t$ [^6] For clarity of argument, in the baseline model I assume all coordinates of $X_{it}$ to be endogenous in the sense that they can be correlated with $\beta_{it}$ through both $A_{i}\text{ and }\eta_{it}$. In fact, the proposed method in this paper can benefit from additional exogenous regressors in (2.1). I discuss it in relation to Assumption 4 and note the changes needed when additional exogenous regressors are included in (2.1) at the end of the proof of Theorem 1. $Y_{it}\in\mathbb{R}$ denotes the scalar-valued outcome variable, and $Z_{it}\in\mathbb{R}^{d_{Z}}$ represents a vector of IVs.
  • $g\left(\cdot\right)$ is an unknown vector-valued function of $Z_{it},\ A_{i},\ \text{and }\eta_{it}$ that determines the value of each coordinate of $X_{it}$.
[…]
To clarify the model’s information structure, first nature draws $\left(A_{i},\varepsilon_{it}\right)$ for each agent and period in the economy. Then, agent $i$ in period $t$ chooses her endogenous $X_{it}$ by solving an optimization problem (e.g., profit maximization) after observing $A_{i}$, some signal $\eta_{it}$ about $\varepsilon_{it}$, and exogenously determined IVs $Z_{it}$, leading to (2.2). Finally, agent $i$ obtains its outcome $Y_{it}$ via (2.1). As an analyst, I observe $\left\{Y_{it},X_{it},Z_{it}\right\}$ for $i=1,...,n$ and $t=1,...,T$ and aim to identify $\mathbb{E}\beta_{it}$ (APE) and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ (LAR). I illustrate the information structure in Figure 1.
Nature: draw ($A_{i}$,$\varepsilon_{it}$)Agent: choose $X_{it}$ given ($Z_{it}$,$A_{i}$,$\eta_{it}$)Agent: realize $Y_{it}$ via (2.1)Analyst: observe ($Y_{it}$,$X_{it}$,$Z_{it}$)
Figure 1: Information Structure
Model (2.1)–(2.2) appears in various economic applications. I describe three of them here.
Example 1 (Heterogeneous Production Function Estimation).
@@
Suppose firm $i$ in year $t$ has a heterogeneous Cobb-Douglas production function in the form of (2.1). The capital and labor elasticities $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)\in\mathbb{R}^{2}$ are modeled as a two-dimensional function of the firm fixed effect $A_{i}$ (e.g., manager ability) and a time-varying random shock $\varepsilon_{it}$ (e.g., realized per-period technology shock). When deciding its input values $X_{it}\in\mathbb{R}^{2}$ for capital and labor, assume firm $i$ knows $A_{i}$ and $\eta_{it}$ (e.g., expected per-period technology shock). Clearly, $\eta_{it}$ is correlated with $A_{i}$ and $\varepsilon_{it}$. Assume firm $i$ also observes some exogenously determined input prices $Z_{it}$, such as interest rates and wages, and that the cost function is given by $c\left(x,z\right)$. Then, firm $i$ chooses input values $X_{it}$ by maximizing its expected profit with the knowledge of $\left(Z_{it},A_{i},\eta_{it}\right)$, i.e.,
$X_{it}=\arg\max_{x\in\mathbb{R}^{2}}\left[\mathbb{E}\left[\left.x^{{}^{\prime}}\beta\left(A_{i},\varepsilon_{it}\right)\right|Z_{it},A_{i},\eta_{it}\right]-c\left(x,Z_{it}\right)\right],$
@@
Example 2 (Labor Supply Estimation).
@@
Suppose agent $i$ has a linear labor supply function of the form (2.1), where $Y_{it}$ represents the number of hours worked and $X_{it}$ includes the endogenous hourly wage along with other exogenous demographic variables. The coordinate of $\beta_{it}$ corresponding to the wage is the key object of interest, as it quantifies how labor supply responds to variations in wage rate over time. Given some exogenous IVs $Z_{it}$ (e.g., county minimum wage and non-labor income), individual capability $A_{i}$, and expected health shocks $\eta_{it}$ (e.g., the likelihood of contracting Covid-19 at work), agent $i$ chooses a job with the wage that maximizes her expected utility, resulting in (2.2).
Example 3 (Engel Curve Estimation).
@@
Suppose the budget share of gasoline $Y_{it}$ for household $i$ at time $t$ is a function of gas price and total expenditure represented in (2.1). Here, $\beta_{it}\in\mathbb{R}^{2}$ is modeled as a vector of unknown functions of the household fixed effect $A_{i}$ and a wealth shock $\varepsilon_{it}$, and captures the elasticity of gasoline demand with respect to total expenditure and gas price, respectively. Given $A_{i}$, a signal $\eta_{it}$ about the wealth shock $\varepsilon_{it}$, and an IV $Z_{it}$ (e.g., head of household’s gross income), household $i$ searches for a gas price and plans its total expenditure budget by maximizing its expected utility, leading to (2.2).
In these examples, the time-varying correlation between $X_{it}$ and $\beta_{it}$ highlights the significance and importance of the time-varying endogeneity through the random coefficients, which naturally arises due to agent optimization behavior. Numerous classic papers (Manski (1987), Altonji and Matzkin (2005), Graham and Powell (2012), Chernozhukov, Fernández-Val, Hahn, and Newey (2013)) have addressed such endogeneity by allowing for the correlation between $X_{it}$ and $A_{i}$. For instance, in production function estimation, a firm with a good management team is typically associated with larger input levels to leverage its managerial skill. This type of correlation has been addressed in the aforementioned papers. However, it carries important empirical implication to allow $X_{it}$ and $\beta_{it}$ to be correlated through factors beyond $A_{i}$. For instance, firms with a good management team may strategically select low levels of inputs for certain periods, which can not be explained only by the correlation between $X_{it}$ and $A_{i}$. To account for such behavior, it is useful to introduce an additional time-varying shock $\eta_{it}$ that contains important information about $\beta_{it}$ and affects the value of $X_{it}$.
Allowing $A_{i}$ to enter both the first-step equation (2.2) and the outcome equation (2.1) is another novel feature of this paper, distinguishing it from classic triangular simultaneous equations models (e.g., Newey, Powell, and Vella (1999), Imbens and Newey (2009)). In these models, the assumption is typically made on the first-step equation that determines the values of regressors that there is only one unobservable scalar-valued random variable $\eta_{it}$ that is correlated with both $A_{i}$ and $\varepsilon_{it}$, i.e., $X_{it}=g\left(Z_{it},\eta_{it}\right)$ instead of $X_{it}=g\left(Z_{it},A_{i},\eta_{it}\right)$. Under regularity conditions, they propose to control for $\eta_{it}$ via either the regression residual from regressing $X_{it}$ on $Z_{it}$ (e.g., Newey, Powell, and Vella (1999)) or the conditional CDF $F_{\left.X_{it}\right|Z_{it}}\left(\left.X_{it}\right|Z_{it}\right)$ which equals $F_{\eta_{it}}\left(\eta_{it}\right)$ (e.g., Imbens and Newey (2009)). However, having $A_{i}$ appear in (2.2) renders the residual and CDF-based methods inapplicable because the one-to-one relationship between the control variable and $\eta_{it}$ no longer holds. Additionally, my model features a nonseparable first-step equation (2.2) in which $A_{i}$ enters $g\left(\cdot\right)$ potentially nonlinearly. The nonseparability and nonlinearity arise naturally from agent optimization behavior as illustrated in the empirical examples. The nonlinearity precludes the use of common demeaning or first-differencing techniques to eliminate $A_{i}$ (e.g., Laage (2024)). Instead, I rely on an index exclusion condition to deal with the fixed effects (e.g., Liu, Poirier, and Shiu (2024)).
@@ 3.1 Identification Idea
The main challenge to identify $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ comes from the correlation between $X_{it}$ and $\beta_{it}$ in model (2.1)–(2.2) through their dependence on the pairwise correlated random variables $A_{i},\ \varepsilon_{it},$and $\eta_{it}$. Suppose I follow the standard identification argument for linear models by taking the conditional expectation of both sides of (2.1) given $X_{it}$:
$\mathbb{E}\left[\left.Y_{it}\right|X_{it}\right]=X_{it}^{{}^{\prime}}\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right].$ (3.1)
Then, because $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ is a function of $X_{it}$, the textbook methods to identify $\mathbb{E}\beta_{it}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ fail.
To see why, first I cannot identify $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ by multiplying both sides of (3.1) by $X_{it}$, taking outer expectation with respect to $X_{it}$, and finally inverting $\mathbb{E}\left[X_{it}X_{it}^{{}^{\prime}}\right]$:
$\mathbb{E}\left[X_{it}Y_{it}\right]=\mathbb{E}\left\{X_{it}X_{it}^{{}^{\prime}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbb{E}}\mathinner{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\left[\left.\beta_{it}\right|X_{it}\right]}\right\},$ (3.2)
since $X_{it}X_{it}^{{}^{\prime}}$ and $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$, a function of $X_{it}$, are not separable inside the expectation on the right hand side (RHS) of (3.2).
Second, the perturbation-based identification method in linear models does not go through either. This method perturbs $X_{it}$ on both sides of (3.1):
$\partial\mathbb{E}\left[\left.Y_{it}\right|X_{it}\right]/\partial X_{it}=\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}X_{it}^{{}^{\prime}}\partial\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]/\partial X_{it}}.$ (3.3)
Again, because $\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]$ is a function of $X_{it}$, by the chain rule there is an additional term color-coded in red showing up on the RHS of (3.3), invalidating the perturbation-based method.
The key identification idea of this paper is to address the time-varying correlation between $X_{it}$ and $\beta_{it}$ using two control variables, $W_{i}$ and $V_{it}$. Conditional on $W_{i}$ and $V_{it}$, the residual variation in $X_{it}$ is determined solely by the exogenous $Z_{it}$. First, since the APE and LAR are generally not identified if the distribution of $\left.A_{i}\right|{\bf X}_{i}$ is left unrestricted (Liu, Poirier, and Shiu (2024)), I exploit an index exclusion condition, $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)\sim_{d}\left.A_{i}\right|W_{i}$, to control for $A_{i}$, where the existence and construction of $W_{i}$ are supported by the established results in the literature (Mundlak (1978), Altonji and Matzkin (2005), Bester and Hansen (2009), Wooldridge (2019), Arkhangelsky and Imbens (2022)). Then, for $\eta_{it}$ I propose a feasible conditional control variable $V_{it}=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right)$ and prove that $V_{it}=F_{\left.\eta_{it}\right|A_{i},W_{i}}\left(\left.\eta_{it}\right|A_{i},W_{i}\right)$, the RHS of which is a one-to-one function of $\eta_{it}$ given $A_{i}$ and $W_{i}$. Since $A_{i}$ is unknown, I apply the LIE to $\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]$ by further conditioning on the unknown $A_{i}$ in the inner conditional expectation to fix $\eta_{it}$ via $V_{it}$, such that conditional on $A_{i},\ V_{it},\text{ and }W_{i}$, the residual variation in $X_{it}$ is entirely driven by the exogenous $Z_{it}$. This allows me to exclude $X_{it}$ from the inner conditional expectation. Next, I exclude $X_{it}$ from the conditioning set of the outer conditional expectation using the index exclusion condition of $W_{i}$ for $A_{i}$, i.e., $\left.A_{i}\right|\left(X_{it},Z_{it},W_{i}\right)\sim_{d}\left.A_{i}\right|W_{i}$. Hence, I have
$\mathbb{E}\left[\left.\beta_{it}\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[\left.\mathbb{E}\left[\left.\beta_{it}\right|X_{it},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}A_{i}},V_{it},W_{i}\right]\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right],$ (3.4)
which by (2.1) implies
$\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]=X_{it}^{{}^{\prime}}\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right].$ (3.5)
@@
Given sufficient variation in $X_{it}$ conditional on $V_{it}$ and $W_{i}$, I identify $\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]$ by either
$\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]=\mathbb{E}\left[\left.X_{it}X_{it}^{{}^{\prime}}\right|V_{it},W_{i}\right]^{-1}\mathbb{E}\left[\left.X_{it}Y_{it}\right|V_{it},W_{i}\right]$ (3.6)
or
$\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]=\partial\mathbb{E}\left[\left.Y_{it}\right|X_{it},V_{it},W_{i}\right]/\partial X_{it}.$ (3.7)
Consequently, I use the LIE to identify
@@
$\mathbb{E}\beta_{it}=\mathbb{E}\left[\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]\right]\text{ and }\mathbb{E}\left[\left.\beta_{it}\right|X_{it}\right]=\mathbb{E}\left[\left.\mathbb{E}\left[\left.\beta_{it}\right|V_{it},W_{i}\right]\right|X_{it}\right].$ (3.8)
3.2 Assumptions and Main Identification Result

### v3

@@ 2 Model
Consider the following baseline TERC model
$Y_{it}$ $=X_{it}^{\prime}\beta\left(A_{i},\varepsilon_{it}\right),$ (2.1)
@@
where $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)\in\mathbb{R}^{d_{X}}$, the central object of interest, is a vector of random coefficients modeled as an unknown vector-valued function $\beta$ of $(A_{i},\varepsilon_{it})$. Here, $A_{i}$ is a fixed effect and $\varepsilon_{it}$ governs the time-varying behavior of $\beta_{it}$; both may have arbitrary dimension. The mapping $\beta$ itself can be time-varying, because $\varepsilon_{it}$ captures time variation in the functional form. $\eta_{it}\in\mathbb{R}^{d_{\eta}}$ is a continuously distributed, time-varying random vector that may be correlated with $A_{i}$ and $\varepsilon_{it}$; its coordinates may also be mutually correlated. This correlation can be important in applications; see the examples below. Let $X_{it}\in\mathbb{R}^{d_{X}}$ denote choice variables (e.g., capital and labor), $Y_{it}\in\mathbb{R}$ the outcome, and $Z_{it}\in\mathbb{R}^{d_{Z}}$ instruments. Finally, $g$ is an unknown vector-valued function of $(Z_{it},A_{i},\eta_{it})$ that determines each coordinate of $X_{it}$.
The TERC model is widely used in empirical studies. It extends the classic constant-coefficient linear panel model while preserving directly interpretable targets—such as the APE and LAR—that carry clear policy relevance. It also accommodates natural extensions, including higher-order moments of the APE and LAR. In what follows, I describe three applications of the TERC model.
Example 1 (Production Function Estimation).
@@
Suppose firm $i$ in year $t$ has a heterogeneous Cobb-Douglas production function in the form of (2.1). The capital and labor elasticities $\beta_{it}=\beta\left(A_{i},\varepsilon_{it}\right)\in\mathbb{R}^{2}$ are modeled as a two-dimensional function of the firm fixed effect $A_{i}$ (e.g., manager ability that is assumed to be stable over time) and a time-varying random shock $\varepsilon_{it}$ (e.g., realized per-period technology shock). When deciding its $X_{it}\in\mathbb{R}^{2}$ of capital and labor, assume firm $i$ knows $A_{i}$ and $\eta_{it}$ (e.g., expected per-period technology shock). Clearly, $\eta_{it}$ is correlated with $A_{i}$ and $\varepsilon_{it}$. Assume firm $i$ also observes $Z_{it}$, which captures local input market dynamics that affect $X_{it}$—for example, competitors’ debt-weighted interest rates and employment-weighted wages. Let the cost function be given by $c\left(x,z\right)$. Then, firm $i$ selects $X_{it}$ by maximizing its expected profit with the knowledge of $\left(Z_{it},A_{i},\eta_{it}\right)$, i.e.,
$X_{it}=\arg\max_{x\in\mathbb{R}^{2}}\left[\mathbb{E}\left[\left.x^{\prime}\beta\left(A_{i},\varepsilon_{it}\right)\right|Z_{it},A_{i},\eta_{it}\right]-c\left(x,Z_{it}\right)\right],$
@@
Example 2 (Labor Supply).
@@
Suppose worker $i$ has a linear labor supply function of the form (2.1), where $Y_{it}$ is the number of hours worked and $X_{it}$ includes the endogenous hourly wage ($X_{it,1}$) along with other exogenous demographic variables. The coordinate of $\beta_{it}$ corresponding to the wage is the key object of interest because it quantifies the elasticity of labor supply with respect to the wage rate. Given exogenous factors $Z_{it}$ (e.g., county minimum-wage changes or earned income tax credit expansions), individual ability $A_{i}$, and shock $\eta_{it}$ (e.g., her expectation of a near-term surge in labor demand in her sector, which increases the expected marginal product of labor and wage offers), she chooses a job with the wage that maximizes her expected utility, resulting in (2.2). Finally, worker $i$ commits $Y_{it}$ hours of working via (2.1).
Example 3 (Almost Ideal Demand System (AIDS)).
@@
Suppose household $i$’s gasoline budget share $Y_{it}$ at time $t$ depends on the gasoline price $X_{it,1}$ and total expenditure $X_{it,2}$ as in (2.1). Let $\beta_{it}\in\mathbb{R}^{2}$ denote the possibly heterogeneous elasticities with respect to $X_{it}$, modeled as unknown functions of a household fixed effect $A_{i}$ (e.g., a persistent propensity to drive versus use transit) and a time-varying wealth shock $\varepsilon_{it}$. Given instruments $Z_{it}$ (e.g., head of household’s earned income and gas tax change), $A_{i}$, and a signal $\eta_{it}$ about $\beta_{it}$ (e.g., expected near-term income change such as upcoming bonus), the household searches for gasoline prices and sets its total expenditure budget by maximizing expected utility, which yields (2.2).
The time-varying correlation between $X_{it}$ and $\beta_{it}$ illustrates a distinct source of endogeneity: optimization-induced selection into regressors through random coefficients. Classic work (Manski (1987), Altonji and Matzkin (2005), Graham and Powell (2012), Chernozhukov, Fernández-Val, Hahn, and Newey (2013)) allows $X_{it}$ to correlate with $A_{i}$ (e.g., better-managed firms choose higher inputs). But empirically, $X_{it}$ can also co-move with $\beta_{it}$ beyond $A_{i}$, because time variation in choices is rarely fully spanned by instrument variation $Z_{it}$. For example, even with little movement in input prices, a well-managed firm may temporarily scale inputs up or down in response to private information about returns. This motivates a time-varying shock $\eta_{it}$ that affects $X_{it}$ and is informative about $\beta_{it}$.
A second departure from standard triangular models (Newey, Powell, and Vella (1999), Imbens and Newey (2009)) is that $A_{i}$ enters both the outcome equation and the first-step choice equation: $X_{it}=g\left(Z_{it},A_{i},\eta_{it}\right)$, not $X_{it}=g\left(Z_{it},\eta_{it}\right)$. Consequently, residual- or CDF-based control-function approaches that rely on a one-to-one mapping between the control and $\eta_{it}$ no longer apply. Moreover, $g$ is nonseparable and can depend on $A_{i}$ nonlinearly, as implied by optimization, so demeaning or first-differencing cannot eliminate $A_{i}$. I instead use an index exclusion condition to handle the fixed effects.
@@ 3.1 Identification Idea
Why Standard Linear Identification Fails
The main obstacle to identifying the APE and the LAR function in the TERC model using standard linear identification is that the regressor is chosen by the agent using information that is also informative about the random coefficients. To see where the standard argument breaks down, take conditional expectations in (2.1) given $X_{it}$:
$\mathbb{E}[\left.Y_{it}\right|X_{it}]=X_{it}^{\prime}\mathbb{E}[\left.\beta_{it}\right|X_{it}].$ (3.1)
In a constant-coefficient model, the term $\mathbb{E}[\left.\beta_{it}\right|X_{it}]$ would be a constant vector and could be recovered from moment conditions. Here, however, $\mathbb{E}[\left.\beta_{it}\right|X_{it}]$ is generally a function of $X_{it}$. As a result, the familiar argument based on multiplying both sides of (3.1) by $X_{it}$, taking expectations, and inverting $\mathbb{E}[X_{it}X_{it}^{\prime}]$ does not identify $\mathbb{E}[\left.\beta_{it}\right|X_{it}]$:
$\mathbb{E}[X_{it}Y_{it}]=\mathbb{E}\left[X_{it}X_{it}^{\prime}\mathbb{E}[\left.\beta_{it}\right|X_{it}]\right],$ (3.2)
because $X_{it}X_{it}^{\prime}$ and $\mathbb{E}[\beta_{it}\mid X_{it}]$ move together inside the expectation. Likewise, the “derivative/perturbation” argument for linear models does not isolate $\mathbb{E}[\left.\beta_{it}\right|X_{it}]$, since differentiating (3.1) yields
$\frac{\partial\mathbb{E}[\left.Y_{it}\right|X_{it}]}{\partial X_{it}}=\mathbb{E}[\left.\beta_{it}\right|X_{it}]+\left(\frac{\partial\mathbb{E}[\left.\beta_{it}\right|X_{it}]}{\partial X_{it}}\right)^{\prime}X_{it},$ (3.3)
and the second term captures how selection into different $X_{it}$ changes the conditional distribution of $\beta_{it}$.
Motivation and Interpretation of the Two Control Variables
The identification strategy in this paper is to recreate, using observables, the thought experiment “vary $X_{it}$ while holding fixed everything the agent knows that is related to $\beta_{it}$.” Doing so requires two controls—one for the time-invariant heterogeneity $A_{i}$ and one for the time-varying information $\eta_{it}$.
The first control is a statistic $W_{i}=W({\bf X}_{i},{\bf Z}_{i})$ constructed from the individual’s panel history. The maintained index exclusion condition implies that once I condition on $W_{i}$, the remaining variation in $(X_{it},Z_{it})$ for a fixed $t$ contains no additional information about $A_{i}$:
$\left.A_{i}\right|(X_{it},Z_{it},W_{i})\sim_{d}\left.A_{i}\right|W_{i}.$ (3.4)
Intuitively, $W_{i}$ is chosen to summarize the persistent, individual-specific component of behavior that is linked to $A_{i}$. For example, in many panel applications a time average of $X_{it}$ (possibly together with time averages of instruments) is a natural candidate: firms with higher managerial ability or individuals with higher baseline productivity tend to exhibit persistently different choices, and $W_{i}$ is designed to capture that persistent component.
Even after accounting for $A_{i}$, endogeneity remains because agents may respond to time-varying information $\eta_{it}$ when choosing $X_{it}$. The second control is constructed as the conditional CDF (a “rank” or “generalized residual”):
$V_{it}:=F_{\left.X_{it}\right|Z_{it},W_{i}}\left(\left.X_{it}\right|Z_{it},W_{i}\right).$ (3.5)
The interpretation is that $V_{it}$ records where the realized $X_{it}$ sits in the distribution of $X_{it}$ among agents with the same $(Z_{it},W_{i})$. Since $A_{i}$ and $Z_{it}$ appear in the first-step equation (1.2) and $W_{i}$ controls for $A_{i}$ by (3.4), the residual variation of $X_{it}$ given $(Z_{it},W_{i})$ is only driven by $\eta_{it}$ which is assumed to enter (1.2) strictly monotonically. For example, consider two firms with the same observed average inputs over all periods ($W_{i}$) and input prices in period $t$ ($Z_{it}$). If firm 1’s choice of capital ($X_{it}$) in this period is higher than that of firm 2, it can be inferred that firm 1 receives a larger productivity shock ($\eta_{it}$) to its capital choice function in period $t$ than firm 2 since (3.4) ensures that the residual variation in $A_{i}$ given $W_{i}$ is not informative about the choice of $X_{it}$.
As a result, $V_{it}$ records where the realized $\eta_{it}$ sits in the distribution of $\eta_{it}$ among agents with the same $(Z_{it},A_{i},W_{i})$. By an exogeneity assumption between $Z_{it}$ and $\eta_{it}$, it drops out from the conditioning set and hence $V_{it}$ controls for $\eta_{it}$ given $(A_{i},W_{i})$. Thus, conditioning on $(V_{it},W_{i})$ is intended to hold fixed both the individual’s persistent type (through $W_{i}$) and the individual’s period-$t$ information that drives endogenous adjustment (through $V_{it}$).
This two-control structure matches the economics of the motivating examples in Section 2. In the production function estimation example, $W_{i}$ captures persistent firm heterogeneity (e.g., management skill) that affects both average input choices and average elasticities, while $V_{it}$ captures time-$t$ private information or expectations about technology that shift input choices and are correlated with realized productivity shocks. In the labor supply application, $W_{i}$ summarizes persistent ability/tastes and $V_{it}$ captures time-varying job-match information that affects wage choice. In the AIDS application, $W_{i}$ captures persistent household heterogeneity and $V_{it}$ captures time-varying wealth or liquidity information that affects expenditure plans.
Steps to Identify APE and LAR
Given the two controls $V_{it}$ and $W_{i}$, the central step to identification of the APE and LAR is to show that once I condition on the two controls, $X_{it}$ no longer carries additional information about $\beta_{it}$. The argument proceeds in two layers.
The first step tackles the time-varying part of the endogeneity by assuming $A_{i}$ is known and showing
$\mathbb{E}[\left.\beta_{it}\right|X_{it},A_{i},V_{it},W_{i}]=\mathbb{E}[\left.\beta_{it}\right|A_{i},V_{it},W_{i}].$ (3.6)
Specifically, since $V_{it}$ records where the realized $\eta_{it}$ sits in the distribution of $\eta_{it}$ among agents with the same $(A_{i},W_{i})$, holding $(A_{i},V_{it},W_{i})$ fixed amounts to holding $(A_{i},\eta_{it})$ fixed. In this step, Assumption 1 (Componentwise Monotonicity) is used to invert $\eta_{it}$ out from function $g$, Assumption 2 (Index Exclusion) allows one to add $A_{i}$ into the conditioning set of the CDF $V_{it}$, and Assumption 3(a) (Exogeneity of IV) enables one to exclude $Z_{it}$ from the conditioning set of the CDF $V_{it}$. Assumption 3(b) (Monotonicity of the Conditional CDF) then implies that $V_{it}$ is a one-to-one transformation of $\eta_{it}$ given $(A_{i},W_{i})$, proving the claim that $(A_{i},V_{it},W_{i})$ uniquely pins down $(A_{i},\eta_{it},W_{i})$. Hence, equation (1.2) implies that the remaining variation in $X_{it}$ is generated only by $Z_{it}$. Under the maintained IV exogeneity condition, this residual variation is independent of $\beta_{it}$. This is the point at which the time-varying part of the endogeneity problem is resolved: within cells defined by $(A_{i},V_{it},W_{i})$, movements in $X_{it}$ are “as good as randomized” by $Z_{it}$. In the production function estimation example, if two firms have the same management ability ($A_{i}$) and expected per-period technology shock ($\eta_{it}$) but selects different capital or labor, such difference can only be explained by the variation in exogenous $Z_{it}$, which is causal and thus can be exploited to identify the APE and LAR.
The second step tackles the time-invariant part of the endogeneity. Since $A_{i}$ is not observed, it is not feasible to condition on it in (3.6). I deal with unknown $A_{i}$ by integrating it out. This involves iterating expectations and the index exclusion restriction of $W_{i}$. Specifically, I show
$\mathbb{E}[\left.\mathbb{E}[\left.\beta_{it}\right|A_{i},V_{it},W_{i}]\right|X_{it},V_{it},W_{i}]=\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}].$ (3.7)
Notice that on the left-hand side of (3.7), $X_{it}$ affects the conditional expectation only through the conditional density $f_{\left.A_{i}\right|X_{it},V_{it},W_{i}}$. The index exclusion restriction (Assumption 2) implies that, conditional on $W_{i}$, $A_{i}$ is independent of $(X_{it},Z_{it})$ and thus of $V_{it}$. For instance, in the production function estimation example, $W_{i}$ captures the firm’s time-invariant management ability $A_{i}$, so once $W_{i}$ is held fixed, period-$t$ inputs $X_{it}$ and control for productivity shock $V_{it}$ only reflect transitory variation and add no extra information about $A_{i}$. Therefore, I have
$f_{\left.A_{i}\right|X_{it},V_{it},W_{i}}(\left.A_{i}\right|X_{it},V_{it},W_{i})=f_{\left.A_{i}\right|W_{i}}(\left.A_{i}\right|W_{i}),$ (3.8)
which ensures that (3.7) holds.
Combining (3.6) and (3.7) gives
$\mathbb{E}[\left.\beta_{it}\right|X_{it},V_{it},W_{i}]=\mathbb{E}\left[\left.\mathbb{E}[\left.\beta_{it}\right|X_{it},A_{i},V_{it},W_{i}]\right|X_{it},V_{it},W_{i}\right]=\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}].$ (3.9)
Equation (3.9) is the key identification result: after conditioning on $(V_{it},W_{i})$, the dependence of $\beta_{it}$ on $X_{it}$ is removed.
Given (3.9), equation (2.1) implies a conditional linear representation
$\mathbb{E}[\left.Y_{it}\right|X_{it},V_{it},W_{i}]=X_{it}^{\prime}\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}].$ (3.10)
@@
With sufficient residual variation in $X_{it}$ conditional on $(V_{it},W_{i})$, $\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}]$ can be identified either from conditional second moments
$\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}]=\mathbb{E}[\left.X_{it}X_{it}^{\prime}\right|V_{it},W_{i}]^{-1}\mathbb{E}[\left.X_{it}Y_{it}\right|V_{it},W_{i}],$ (3.11)
or from the derivative of the conditional mean
$\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}]=\frac{\partial\mathbb{E}[\left.Y_{it}\right|X_{it},V_{it},W_{i}]}{\partial X_{it}}.$ (3.12)
Finally, identification of the APE and LAR follows by the law of iterated expectations
@@
$\overline{b}=\dfrac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}]\right],\quad\text{and}\quad b_{t}(X_{it})=\mathbb{E}\left[\left.\mathbb{E}[\left.\beta_{it}\right|V_{it},W_{i}]\right|X_{it}\right].$ (3.13)
In the production function estimation example, $\overline{b}$ summarizes output elasticities by averaging across firms and time, so it plays the same role as the constant coefficient in standard linear panel regressions. $b_{t}(x)$ denotes the period-$t$ average elasticity conditional on operating at input level $x$, a policy-relevant local parameter when elasticities are systematically related to firms’ input choices. Analogous interpretations apply to labor supply (average and local wage elasticities) and demand estimation (average and local expenditure/price elasticities). The purpose of $(W_{i},V_{it})$ is precisely to make these interpretations credible by ensuring that the remaining variation used to identify (3.13) is driven by the exogenous instruments rather than by unobserved, coefficient-relevant information.
3.2 Assumptions